During a thermodynamic process, the pressure P of a diatomic gas depends on its volume V as P ∝ V^-4 . The specific heat of gas during this process

12R

13R/6

17R/5

11R/2

Grade:12

1 Answers

Vikas TU
8 years ago
This is the case of Polytropic process and therefore,
C is given as:
C = R/(Y -1) + R/(1-x)
where Y = Cp/Cv.
where x = 4 given.
Cp =( F/2+1)*R
Cv = FR/2
F = 5 for diaatomic gas that is degree of freedom.